# SafePTR[Uncaptioned image]: Token-Level Jailbreak Defense in Multimodal LLMs via Prune-then-Restore Mechanism

### v1

@@ Abstract
By incorporating visual inputs, Multimodal Large Language Models (MLLMs) extend LLMs to support visual reasoning. However, this integration also introduces new vulnerabilities, making MLLMs susceptible to multimodal jailbreak attacks and hindering their safe deployment. Existing defense methods, including Image-to-Text Translation, Safe Prompting, and Multimodal Safety Tuning, attempt to address this by aligning multimodal inputs with LLMs’ built-in safeguards. Yet, they fall short in uncovering root causes of multimodal vulnerabilities, particularly how harmful multimodal tokens trigger jailbreak in MLLMs? Consequently, they remain vulnerable to text-driven multimodal jailbreaks, often exhibiting overdefensive behaviors and imposing heavy training overhead. To bridge this gap, we present an comprehensive analysis of where, how and which harmful multimodal tokens bypass safeguards in MLLMs. Surprisingly, we find that less than 1% tokens in early-middle layers are responsible for inducing unsafe behaviors, highlighting the potential of precisely removing a small subset of harmful tokens, without requiring safety tuning, can still effectively improve safety against jailbreaks. Motivated by this, we propose Safe Prune-then-Restore (SafePTR), an training-free defense framework that selectively prunes harmful tokens at vulnerable layers while restoring benign features at subsequent layers. Without incurring additional computational overhead, SafePTR significantly enhances the safety of MLLMs while preserving efficiency. Extensive evaluations across three MLLMs and five benchmarks demonstrate SafePTR’s state-of-the-art performance in mitigating jailbreak risks without compromising utility.
@@ 1 Introduction
Multimodal large language models (MLLMs)OpenAI (2023); Lu et al. (2024); Liu et al. (2023a); Bai et al. (2023); Zhu et al. (2023) extend the capabilities of large language models (LLMs)Touvron et al. (2023); Chiang et al. (2023); Jiang et al. (2024) to visual inputs, enabling unified language–vision reasoning. Despite strong performance across multimodal tasks, MLLMs remain vulnerable to multimodal jailbreak attacks Luo et al. (2024b); Gong et al. (2025); Dong et al. (2023), raising critical safety concerns for secure deployment.
@@
Figure 1: (Left) Existing MLLM defense methods remain susceptible to text-driven multimodal jailbreaks, exhibiting overdefensive behavior and imposing heavy training overhead. (Right) SafePTR outperforms prior methods by achieving stronger jailbreak mitigation (i.e., Jailbreak28K, Figstep and MM-Safety), better preserving task utility (i.e., MMVet and MME), and minimal computational overhead (i.e., Training-free and One-bypass Inference). Performances of SafePTR across more MLLMs are provided in Appendix.A
@@
To mitigate safety risks in MLLMs, existing approaches Gou et al. (2024); Guo et al. (2024); Ghosal et al. (2024); Gao et al. (2024); Gong et al. (2025) primarily adapt multimodal inputs to align with the built-in safeguard mechanisms of underlying LLMs, inherited from prior safety alignment processes conducted during LLMs’ safety fine-tuning stage. These methods generally fall into three categories: (1) Image-to-Text Translation Gou et al. (2024); Guo et al. (2024), which converts visual inputs into textual descriptions to leverage LLM’s existing safety boundaries, but remains susceptible to text-driven multimodal jailbreaks (Fig.1 (a)); (2) Safe Prompting Wang et al. (2024b); Gong et al. (2025), which injects static safety constraints into instructions to suppress unsafe completions, but often leads to overdefensive behavior that degrade model’s utility on benign inputs (Fig.1 (b)); and (3) Multimodal Safety Tuning Liu et al. (2024b, c) leverages safety-specific datasets to either train dedicated risk detectors Ghosal et al. (2024) or align multimodal representations with LLM safety priors via cross-modal supervision Gao et al. (2024) While effective, these methods incur substantial training overhead and exhibit limited generalization to unseen jailbreaks (see Fig. 1(c)).
@@
The root cause lies in existing methods’ reliance on the built-in safeguards of LLMs without uncovering the underlying mechanisms of multimodal vulnerabilities, particularly “how harmful multimodal tokens exploit internal representation pathways to trigger jailbreaks?”, thereby limiting the development of interpretable defenses tailored for MLLMs. Specifically, (1) Image-to-Text Translation ignores where vulnerabilities occur by bypassing visual processing layers. By directly converting images into text and relying solely on LLM’s built-in safeguards, it remains susceptible to text-driven multimodal jailbreaks, particularly when the adversarial textual prompts (e.g., from JailbreakV-28K Luo et al. (2024b)) already evade these defenses; (2) Safe Prompting uniformly applies static safety constraints across various inputs, without adaptively modeling how unsafe behaviors emerge in different contexts; This lack of adaptivity prevents the model from identifying fine-grained safety concepts, such as distinguishing “a toy water gun” from “a real weapon” in Fig. 1(b), often resulting in overdefensive responses that compromise utility in benign scenarios; (3) Multimodal Safety Tuning enforces safety alignment based on dataset-specific defense preferences, but fails to identify which multimodal tokens trigger jailbreaks, limiting interpretability and adaptability. For example, TGA Liu et al. (2024c) relies on 1223K samples to train a safety preference model, incurring high cost (involving 64 × V100 GPUs) while offering limited robustness across diverse jailbreak scenarios.
@@ 2 Harmful Token Propagation and Jailbreak Activation
Experimental Setting: We study three MLLMs, LLaVA-1.5-7B Liu et al. (2023a), MiniGPT-4 Zhu et al. (2023), and DeepSeek-VL Lu et al. (2024), which exhibit notable jailbreak vulnerabilities. To analyze how malicious semantics propagate, we use two multimodal jailbreak datasets: (1) FigStep Gong et al. (2023), which transforms harmful instructions into typographic images across 10 prohibited categories (500 text-image pairs); and (2) MM-SafetyBench Liu et al. (2023b), which employs jailbreak images from Stable Diffusion and typography across 13 restricted scenarios (5,040 pairs). Once processed by the MLLMs, these images are encoded into tokens whose harmful semantics influence model behavior across layers.
Moreover, following (Wang et al., 2024b; Gao et al., 2024; Luo et al., 2024a), we adopt the Attack Success Rate (ASR) to measure the vulnerability of these models to multimodal jailbreak attacks. Specifically, given a test dataset $\mathcal{D}_{\text{unsafe}}$ of crafted jailbreak image-text pairs, the ASR quantifies the ratio of harmful responses to the total number of input queries and is defined as:
@@
Finding-1 (where): A few early-middle layers are especially vulnerable to harmful tokens. We conduct a Layer-wise Intervention Analysis (LIA) (see Appendix.B for details) to identify layers most susceptible to malicious inputs by sequentially removing hidden states of the attack-triggering modality (visual or textual) and observing their impact on model behavior. Specifically, for each layer window $[n,n+\Delta_{n}]$, we disable the influence of the malicious modality. As shown in Fig. 2, the resulting changes in Attack Success Rate (ASR) across layers reveal the model’s vulnerability profile.
Figure 2: Layer-wise vulnerability analysis of MLLMs. Each curve represents the Attack Success Rate (ASR) under layer-wise interventions with varying contiguous layer spans $k={2,4}$ The orange region highlights the layers most susceptible to safety breaches, with its left and right boundaries marking the earliest and latest compromised layers within the model, respectively. Since the intervention requires $k$ consecutive layers, the horizontal axis is limited to the range $[0,L-k]$.
Through Layer-wise vulnerability analysis, we observe that while harmful tokens propagate through all layers, their impact on attack success varies significantly. For LLaVA-1.5-7B, DeepSeek-VL2, and MiniGPT-4-7B, pruning harmful tokens in just 2–4 consecutive early–middle layers (e.g., $[7,9)$, $[4,6)$, or $[7,9)$) significantly reduces ASR from 67.3% to 4.2%, revealing that jailbreak attacks mainly exploit a narrow band of contiguous vulnerable layers. In contrast, pruning subsequent “safety layers” provides limited defensive benefit, as they are primarily responsible for cross-modal integration and language refinement Yue et al. (2024); Liu et al. (2024a). This is further supported by our ablation study 4.4, which shows that restoring benign features in these layers significantly enhances utility.
Finding-2 (how): Greater semantic deviation from safety alignment increases jailbreak susceptibility. Building on Finding-1, we further examine how jailbreaks manifest in layers most susceptible to harmful tokens’ influence. Considering that the safety-aligned instructions in AdaShield Wang et al. (2024b) have been shown to effectively enhance model safety, we adopt them as semantic references to define the model’s aligned safety space. Specifically, we compute both Cosine similarity (y-axis) and Euclidean distance (x-axis) between the hidden states of input samples and those within safety-aligned instructions.
@@
Table 1: Harmful Token Ratio (%) across LLaVA-1.5, MiniGPT-4 and DeepSeek-VL2 on Figstep and MM-SafetyBench.
Finding-3 (which): A small fraction of harmful tokens induce significant semantic deviation from safety-aligned tokens. Building on Finding-2, we further investigate which individual tokens are responsible for such semantic deviation. To this end, we compute the semantic distance($S^{l}=1-cos(V^{l},R^{l})$) between tokens ($V^{l}$) and the safety-aligned instruction representation ($R^{l}$) at layer-$l$ across samples from both FigStep (500) and MM-SafetyBench (5040). Tokens whose deviation($s_{ij}^{l}$) exceeds the threshold ($\alpha$) are marked as harmful ones ($v_{ij}^{l}\in V^{l}$). As shown in Fig. 4, for LLaVA-1.5-7B, an example from MM-SafetyBench illustrates that harmful tokens are sparsely distributed and primarily concentrated in early-middle layers (e.g., [7, 9)), accounting for only 0.62% of all input tokens. Similar patterns are observed on Figstep across MiniGPT-4-7B (0.93%) and DeepSeek-VL2 (1.66%), as summarized in Tab. 1. This behavior may stem from attention sinks, where certain tokens receive excessive attention and dominate the model’s internal representations, as discussed in prior studies Ma et al. (2023); Zhang et al. (2024b). To further analyze their impact, we visualize both a semantic deviation heatmap ($S^{8}$) and an overlaid binary mask ($V^{\prime 8}=\{v_{ij}^{8}\in V^{8}\}$) at layer 8 using a sample from MM-SafetyBench. Tokens representing “armed figures,” “smoke,” and “terrain” associated with violent scenarios exhibit high semantic drift. Interestingly, several background tokens also show large deviations, suggesting that semantically deviant cues from inconspicuous regions can disrupt overall structure, amplify misalignment, and weaken model’s safeguards against jailbreaks.
@@ 3.1 Harmful Token Prunning
As indicated in Fig. 2, the very early layers, contribute minimally to harmful behaviors For the subsequent layers $[n,n+\Delta_{n})$, where $\Delta_{n}=2$ (e.g., layers $[7,9)$ for LLaVA1.5-7B and MiniGPT-4-7B, layers $[4,6)$ for DeepSeek-VL2), Harmful Token Prunning (HTP) is applied to remove vision/instruction harmful tokens, as shown in Eqn. 2.
$[H_{img}^{l+1},H_{ins}^{l+1},H_{safe}^{l+1}]=\textrm{FFN}^{l}(\textrm{Attention}^{l}([HTP(H_{img}^{l},H_{ins}^{l},H_{safe}^{l})]),l\in[n,n+\Delta_{n}].$ (2)
Here, $\textrm{{H}}_{ins}^{l}=\in\mathbb{R}^{T\times D}$, $\textrm{{H}}_{img}^{l}\in\mathbb{R}^{M\times D}$ and $\textrm{{H}}_{safe}^{l}\in\mathbb{R}^{C\times D}$ denotes the hidden states of user instructions, input images and safety-aligned instruction at $l\;th$ layer. Due to variations in harmful content across inputs and differences in alignment strength across different models, a universal threshold for identifying harmful tokens is impractical. Instead, we adopt Top-K selection strategy Chen et al. (2024) and empirically determine a globally effective Top-K value to guide the pruning process, as shown in Tab. 9. By computing the distance between visual tokens and hidden states of safe-aligned instructions, i.e., “Please answer this question safely and accurately. Reply ‘I am sorry.’ if the content is harmful.” in Eqn. 3,
@@ 3.2 Benign Features Restoration
After pruning harmful tokens in vulnerable layers , the subsequent layers operate on incomplete visual representations, lacking fine-grained contextual interactions. To restore utility without compromising robustness, we introduce the Benign Feature Restoration (BFR) module, which selectively reintegrates benign features while avoiding harmful semantics. Formally, the restored hidden states are computed as: $SH_{img}^{n+\Delta_{n}}=BFR(\hat{H}_{img}^{n+\Delta_{n}-1},H_{img}^{n+\Delta_{n}-1})$. Specifically, while HTP prunes harmful tokens within layers $[n,n+\Delta_{n}]$, BFR maintains a parallel branch that performs standard inference over the same layers and selectively restores benign features. This dual-path design enables the recovery of hidden states in subsequent layers, which are less susceptible to attack and primarily responsible for cross-modal integration and language refinement. By restoring previously pruned tokens at this stage, the model preserves functional performance without compromising safety. As shown in Eqn. 4:
$[H_{img}^{l+1},H_{ins}^{l+1},H_{safe}^{l+1}]=\textrm{FFN}^{l}(\textrm{Attention}^{l}([SH_{img}^{l},SH_{ins}^{l},H_{safe}^{l}]),l=n+\Delta_{n}.$ (4)
@@ 4.1 Experimental Details
Implementation Details. Following Immune Ghosal et al. (2024), we implement the proposed SafePTR using Hugging Face Transformers library. The LLaVA1.5-7B results are based on version 1.2.2 from the official benchmark repository. We set the number of tokens sampled k = 10%. For LLaVA-1.5-7B, DeepSeek-VL2, and MiniGPT-4-7B, harmful tokens are pruned in layers $[7,9)$, $[4,6)$, $[7,9)$. All experiments are conducted on four RTX3090 GPUs.
Baseline and Comparable Methods. We evaluate SafePTR on three state-of-the-art open-source MLLMs: LLaVA-1.5-7B (Liu et al., 2023a), MiniGPT-4-7B (Zhu et al., 2023), and DeepSeek-VL2-Tiny (Lu et al., 2024). Comparisons are made against recent jailbreak defense methods, Immune (Wang et al., 2024a), Adashield (Wang et al., 2024b), ECSO (Gou et al., 2024), CoCA (Gao et al., 2024), and FigStep (Gong et al., 2023), under a unified test set and consistent metrics.
Evaluation Benchmarks and Metrics. We assess model performance across three aspects: (1) Safety: We use JailbreakV-28K Luo et al. (2024b) (text-driven), MM-SafetyBench Liu et al. (2023b), and FigStep Gong et al. (2025) (image-driven), reporting Attack Success Rate (ASR) Wang et al. (2024b); Gao et al. (2024). (2) Utility: Benign task accuracy is measured on MME Fu et al. (2023) and MM-Vet Yu et al. (2024), which evaluate multimodal understanding and visual reasoning. (3) Efficiency: We report training data size (K) and inference latency (sec/sample) to evaluate efficiency.
@@
SafePTR 0.0 11.2 0.0     0.0 0.0 0.0     3.6 9.0 9.0     0.0 3.0 1.2     6.2 5.5 11.0     0.6 2.5 0.6     3.5
Table 3: Evaluation on MM-SafetyBench. We report the Attack Success Rate (ASR↓) across six prohibited categories, using GPT-4 as the jailbreak classifier. Bold highlights the best (i.e., lowest) ASR values. The attack image types include typography-based images (T), visuals generated by Stable Diffusion (S), and Stable Diffusion images with overlaid typography subtitles (S-T).

### v2

@@ Abstract
By incorporating visual inputs, Multimodal Large Language Models (MLLMs) extend LLMs to support visual reasoning. However, this integration also introduces new vulnerabilities, making MLLMs susceptible to multimodal jailbreak attacks and hindering their safe deployment. Existing defense methods, including Image-to-Text Translation, Safe Prompting, and Multimodal Safety Tuning, attempt to address this by aligning multimodal inputs with LLMs’ built-in safeguards. Yet, they fall short in uncovering root causes of multimodal vulnerabilities, particularly how harmful multimodal tokens trigger jailbreak in MLLMs? Consequently, they remain vulnerable to text-driven multimodal jailbreaks, often exhibiting overdefensive behaviors and imposing heavy training overhead. To bridge this gap, we present an comprehensive analysis of where, how and which harmful multimodal tokens bypass safeguards in MLLMs. Surprisingly, we find that less than 1% tokens in early-middle layers are responsible for inducing unsafe behaviors, highlighting the potential of precisely removing a small subset of harmful tokens, without requiring safety tuning, can still effectively improve safety against jailbreaks. Motivated by this, we propose Safe Prune-then-Restore (SafePTR), an training-free defense framework that selectively prunes harmful tokens at vulnerable layers while restoring benign features at subsequent layers. Without incurring additional computational overhead, SafePTR significantly enhances the safety of MLLMs while preserving efficiency. Extensive evaluations across three MLLMs and five benchmarks demonstrate SafePTR’s state-of-the-art performance in mitigating jailbreak risks without compromising utility. Our code is available at https://github.com/BT-C/SafePTR.
@@ 1 Introduction
Multimodal large language models (MLLMs)(GPT4V; lu2024deepseek; liu2023visualllava; bai2023qwen; zhu2023minigpt4) extend the capabilities of large language models (LLMs)(touvron2023llama; vicuna2023; jiang2024mixtral) to visual inputs, enabling unified language–vision reasoning. Despite strong performance across multimodal tasks, MLLMs remain vulnerable to multimodal jailbreak attacks (luo2024jailbreakv28k; gong2025figstep; dong2023visual), raising critical safety concerns for secure deployment.
@@
Figure 1: (Left) Existing MLLM defense methods remain susceptible to text-driven multimodal jailbreaks, exhibiting overdefensive behavior and imposing heavy training overhead. (Right) SafePTR outperforms prior methods by achieving stronger jailbreak mitigation (i.e., Jailbreak28K, Figstep and $M$-Safety), better preserving task utility (i.e., MMVet and MME), and minimal computational overhead (i.e., Training-free and One-bypass Inference). Performances of SafePTR across more MLLMs are provided in Appendix.A
@@
To mitigate safety risks in MLLMs, existing approaches (gou2024eyesECSO; guo2024vllmparadox; ghosal2024immune; gao2024coca; gong2025figstep) primarily adapt multimodal inputs to align with the built-in safeguard mechanisms of underlying LLMs, inherited from prior safety alignment processes conducted during LLMs’ safety fine-tuning stage. These methods generally fall into three categories: (1) Image-to-Text Translation (gou2024eyesECSO; guo2024vllmparadox), which converts visual inputs into textual descriptions to leverage LLM’s existing safety boundaries, but remains susceptible to text-driven multimodal jailbreaks (Fig.1 (a)); (2) Safe Prompting (wang2024adashield; gong2025figstep), which injects static safety constraints into instructions to suppress unsafe completions, but often leads to overdefensive behavior that degrade model’s utility on benign inputs (Fig.1 (b)); and (3) Multimodal Safety Tuning (liu2024protector; liu2024crossmodal) leverages safety-specific datasets to either train dedicated risk detectors (ghosal2024immune) or align multimodal representations with LLM safety priors via cross-modal supervision (gao2024coca) While effective, these methods incur substantial training overhead and exhibit limited generalization to unseen jailbreaks (see Fig. 1(c)).
@@
The root cause lies in existing methods’ reliance on the built-in safeguards of LLMs without uncovering the underlying mechanisms of multimodal vulnerabilities, particularly “how harmful multimodal tokens exploit internal representation pathways to trigger jailbreaks?”, thereby limiting the development of interpretable defenses tailored for MLLMs. Specifically, (1) Image-to-Text Translation ignores where vulnerabilities occur by bypassing visual processing layers. By directly converting images into text and relying solely on LLM’s built-in safeguards, it remains susceptible to text-driven multimodal jailbreaks, particularly when the adversarial textual prompts (e.g., from JailbreakV-28K (luo2024jailbreakv28k)) already evade these defenses; (2) Safe Prompting uniformly applies static safety constraints across various inputs, without adaptively modeling how unsafe behaviors emerge in different contexts; This lack of adaptivity prevents the model from identifying fine-grained safety concepts, such as distinguishing “a toy water gun” from “a real weapon” in Fig. 1(b), often resulting in overdefensive responses that compromise utility in benign scenarios; (3) Multimodal Safety Tuning enforces safety alignment based on dataset-specific defense preferences, but fails to identify which multimodal tokens trigger jailbreaks, limiting interpretability and adaptability. For example, TGA (liu2024crossmodal) relies on 1223K samples to train a safety preference model, incurring high cost (involving 64 × V100 GPUs) while offering limited robustness across diverse jailbreak scenarios.
@@ 2 Harmful Token Propagation and Jailbreak Activation
Experimental Setting: We study three MLLMs, LLaVA-1.5-7B (liu2023visualllava), MiniGPT-4 (zhu2023minigpt4), and DeepSeek-VL (lu2024deepseek), which exhibit notable jailbreak vulnerabilities. To analyze how malicious semantics propagate, we use two multimodal jailbreak datasets: (1) FigStep (gong2023figstep), which transforms harmful instructions into typographic images across 10 prohibited categories (500 text-image pairs); and (2) MM-SafetyBench (liu2023mmsafetybench), which employs jailbreak images from Stable Diffusion and typography across 13 restricted scenarios (5,040 pairs). Once processed by the MLLMs, these images are encoded into tokens whose harmful semantics influence model behavior across layers.
Moreover, following (wang2024adashield; gao2024coca; luo2024jailbreakv), we adopt the Attack Success Rate (ASR) to measure the vulnerability of these models to multimodal jailbreak attacks. Specifically, given a test dataset $\mathcal{D}_{\text{unsafe}}$ of crafted jailbreak image-text pairs, the ASR quantifies the ratio of harmful responses to the total number of input queries and is defined as:
@@
Figure 2: Layer-wise vulnerability analysis of MLLMs. Each curve represents the Attack Success Rate (ASR) under layer-wise interventions with varying contiguous layer spans k=2,4k={2,4} . The orange region highlights the layers most susceptible to safety breaches, with its left and right boundaries marking the earliest and latest compromised layers within the model, respectively. Since the intervention requires $k$ consecutive layers, the horizontal axis is limited to the range [0,Lk][0,L-k].
Finding-1 (where): A few early-middle layers are especially vulnerable to harmful tokens. We conduct a Layer-wise Intervention Analysis (LIA) (see Appendix.B for details) to identify layers most susceptible to malicious inputs by sequentially removing hidden states of the attack-triggering modality (visual or textual) and observing their impact on model behavior. Specifically, for each layer window $n,n+\Delta_{n}]$, we disable the influence of the malicious modality. As shown in Fig. 2, the resulting changes in Attack Success Rate (ASR) across layers reveal the model’s vulnerability profile.
Through Layer-wise vulnerability analysis, we observe that while harmful tokens propagate through all layers, their impact on attack success varies significantly. For LLaVA-1.5-7B, DeepSeek-VL2, and MiniGPT-4-7B, pruning harmful tokens in just 2–4 consecutive early–middle layers (e.g., $[7,9)$, $[4,6)$, or $[7,9)$) significantly reduces ASR from 67.3% to 4.2%, revealing that jailbreak attacks mainly exploit a narrow band of contiguous vulnerable layers. In contrast, pruning subsequent “safety layers” provides limited defensive benefit, as they are primarily responsible for cross-modal integration and language refinement (less; eRAM-V). This is further supported by our ablation study 4.4, which shows that restoring benign features in these layers significantly enhances utility.
Finding-2 (how): Greater semantic deviation from safety alignment increases jailbreak susceptibility. Building on Finding-1, we further examine how jailbreaks manifest in layers most susceptible to harmful tokens’ influence. Considering that the safety-aligned instructions in AdaShield (wang2024adashield) have been shown to effectively enhance model safety, we adopt them as semantic references to define the model’s aligned safety space. Specifically, we compute both Cosine similarity (y-axis) and Euclidean distance (x-axis) between the hidden states of input samples and those within safety-aligned instructions.
@@
Table 1: Harmful Token Ratio (%) across LLaVA-1.5, MiniGPT-4 and DeepSeek-VL2 on Figstep and $M$-SafetyBench.
[…]
@@ 3.1 Harmful Token Prunning
As indicated in Fig. 2, the very early layers, contribute minimally to harmful behaviors For the subsequent layers $n,n+\Delta_{n})$, where $\Delta_{n}=2$ (e.g., layers $[7,9)$ for LLaVA1.5-7B and MiniGPT-4-7B, layers $[4,6)$ for DeepSeek-VL2), Harmful Token Prunning (HTP) is applied to remove vision/instruction harmful tokens, as shown in Eqn. 2.
$H_{img}^{l+1},H_{ins}^{l+1},H_{safe}^{l+1}]=\textrm{FFN}^{l}(\textrm{Attention}^{l}([HTP(H_{img}^{l},H_{ins}^{l},H_{safe}^{l})]),l\in[n,n+\Delta_{n}].$ (2)
Here, $\textrm{{H}}_{ins}^{l}=\in\mathbb{R}^{T\times D}$, $\textrm{{H}}_{img}^{l}\in\mathbb{R}^{M\times D}$ and $\textrm{{H}}_{safe}^{l}\in\mathbb{R}^{C\times D}$ denotes the hidden states of user instructions, input images and safety-aligned instruction at lthl\;th layer. Due to variations in harmful content across inputs and differences in alignment strength across different models, a universal threshold for identifying harmful tokens is impractical. Instead, we adopt Top-K selection strategy (chen2024fastv) and empirically determine a globally effective Top-K value to guide the pruning process, as shown in Tab. 9. By computing the distance between visual tokens and hidden states of safe-aligned instructions, i.e., “Please answer this question safely and accurately. Reply ‘I am sorry.’ if the content is harmful.” in Eqn. 3,
@@ 3.2 Benign Features Restoration
After pruning harmful tokens in vulnerable layers , the subsequent layers operate on incomplete visual representations, lacking fine-grained contextual interactions. To restore utility without compromising robustness, we introduce the Benign Feature Restoration (BFR) module, which selectively reintegrates benign features while avoiding harmful semantics. Formally, the restored hidden states are computed as: $SH_{img}^{n+\Delta_{n}}=BFR(\hat{H}_{img}^{n+\Delta_{n}-1},H_{img}^{n+\Delta_{n}-1})$. Specifically, while HTP prunes harmful tokens within layers $n,n+\Delta_{n}]$, BFR maintains a parallel branch that performs standard inference over the same layers and selectively restores benign features. This dual-path design enables the recovery of hidden states in subsequent layers, which are less susceptible to attack and primarily responsible for cross-modal integration and language refinement. By restoring previously pruned tokens at this stage, the model preserves functional performance without compromising safety. As shown in Eqn. 4:
$H_{img}^{l+1},H_{ins}^{l+1},H_{safe}^{l+1}]=\textrm{FFN}^{l}(\textrm{Attention}^{l}([SH_{img}^{l},SH_{ins}^{l},H_{safe}^{l}]),l=n+\Delta_{n}.$ (4)
@@ 4.1 Experimental Details
Implementation Details. Following Immune (ghosal2024immune), we implement the proposed SafePTR using Hugging Face Transformers library. The LLaVA1.5-7B results are based on version 1.2.2 from the official benchmark repository. We set the number of tokens sampled k = 10%. For LLaVA-1.5-7B, DeepSeek-VL2, and MiniGPT-4-7B, harmful tokens are pruned in layers $[7,9)$, $[4,6)$, $[7,9)$. We repeat the experiments five times for each metric with different random seeds. All experiments are conducted on four RTX3090 GPUs.
Baseline and Comparable Methods. We evaluate SafePTR on three state-of-the-art open-source MLLMs: LLaVA-1.5-7B (liu2023visualllava), MiniGPT-4-7B (zhu2023minigpt4), and DeepSeek-VL2-Tiny (lu2024deepseek). Comparisons are made against recent jailbreak defense methods, Immune (wang2024immune), Adashield (wang2024adashield), ECSO (gou2024eyesECSO), CoCA (gao2024coca), and FigStep (gong2023figstep), under a unified test set and consistent metrics.
Evaluation Benchmarks and Metrics. We assess model performance across three aspects: (1) Safety: We use JailbreakV-28K (luo2024jailbreakv28k) (text-driven), MM-SafetyBench (liu2023mmsafetybench), and FigStep (gong2025figstep) (image-driven), reporting Attack Success Rate (ASR) (wang2024adashield; gao2024coca). (2) Utility: Benign task accuracy is measured on MME (fu2023mme) and MM-Vet (yu2024mm), which evaluate multimodal understanding and visual reasoning. (3) Efficiency: We report training data size (K) and inference latency (sec/sample) to evaluate efficiency.
@@
SafePTR 11.7 37.7 36.1     6.9 21.4 20.4     3.6 9.0 9.0     0.0 17.1 13.6     22.5 25.9 28.8     0.6 31.1 32.9     18.2
Table 3: Evaluation on $M$-SafetyBench. We report the Attack Success Rate (ASR↓) across six prohibited categories, using GPT-4 as the jailbreak classifier. Bold highlights the best (i.e., lowest) ASR values. The attack image types include typography-based images (T), visuals generated by Stable Diffusion (S), and Stable Diffusion images with overlaid typography subtitles (S-T).